How do I determine if I use Standard Normal Probabilities or Critical Value. I found the test statistic Z but I am not sure how to figure out if I look for it under the Standard Norm. Probabilities table or the Critical Value table. The questions for each of these look the same to me.

This problem uses critical value table but how is it decided to use that table and not the other?

Using traditional methods, it takes 108 hours to receive a basic flying license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 220 students and observed that they had a mean of 107 hours. Assume the standard deviation is known to be 8. A level of significance of 0.1 will be used to determine if the technique performs differently than the traditional method. Is there sufficient evidence to support the claim that the technique performs differently than the traditional method?
Answer #1

Here, n > 30 so we used z test

Critical value is calculated as:
Here, it is two tailed test

Here, alpha = 0.10

The z value at 90% confidence interval is,

alpha = 1 - 0.90 = 0.1
alpha/2 = 0.1/2 = 0.05
Zalpha/2 = Z0.05 = 1.645


Critical value = 1.645 using z right tailed test

H0 : mu = 108
ha: mu not equasl to 108

z = (xbar - mu)/(s/sqrt(n))
= ( 107 - 108)/98/sqrt(220)
= -0.0007

p value = 2 *P(z< -0.0007)
= 2 * 0.4997
= 0.9995